a bulb is rated 40watt,220volts.it works 10hours a day and is connected to mains.if the mains have 200volts potential differencethen what is its energy consumption for a day?

Grade:10

1 Answers

Ramesh V
14 years ago

power = V2/R

so, resistance of bulb = 220*220/40 = 1210 ohms

now its connected to 200V source

power consumed is P=V2/R = 200*200/1210 = 33W

energy consumption= p*t = 33*10 = 330 W-hr

                                                   = 0.33KW-hr

                  or 0.33 units

--

regards

Ramesh